(x-a)(x-b)
8 ,1
2, 2

Found 3 solutions by Alan3354, jorel555, ikleyn:
Answer by Alan3354(69443)

Do you have a question?

Answer by jorel555(1290)

x-8(x-1)=x^2-9x+8
x-2(x-2)=x^2-4x+4......

Answer by ikleyn(52780)

.
Both the question and the answer are unclear.

Actually, it is the case when the post MUST be ignored/declined since it doesn't contain a question.

With the goal of educating the sender.